Wilson's has 10,000 shares of common stock outstanding at a market price of $35 a share. The firm also has a bond issue outstanding with a total face value of $250,000 which is selling for 102 % of face value. The cost of equity is 11 % while the pre-tax cost of debt is 8 %. The firm has a beta of 1.1 and a tax rate of 34 %. What is Wilson's weighted average cost of capital? 


Question 15 options:

6.50 %


8.72 %


9.08 %


9.63 %


None of the above
Answer #1

Given about Wilson,

share outstanding = 10000

current stock price - $35

So, market value of common stock = shares*price = 10000*35 = $350000

face value of bond = $250000

price/market value of debt = 102% of face value = 102% of 250000 = $255000

So, weight of common stock We = MV of stock/(MV of stock + MV of debt) = 350000/(350000+255000) = 57.85%

Similarly, weight of debt Wd = 1-We = 1-0.5785 = 42.15%

cost of equity Ke = 11%

pretax cost of debt Kd = 8%

tax rate T = 34%

So, WACC = Wd*Kd*(1-T) + We*Ke = 0.4215*8*(1-0.34) + 0.5785*11 = 8.59%

So, Weighted average cost of capital for the firm is 8.59%

So, none of the options are correct.